# Stabilizer Entanglement as a Magic Highway

### v2

@@ Abstract
where $N=|A|+|B|$ is the total number of qubits. This result reveals an intriguing mechanism by which stabilizer entanglement facilitates magic saturation. Notice that the average value of $\overline{Y}$ for Haar random states is given by $4\cdot 2^{-N}$. Eq. (LABEL:eq:UaUb) shows that the gap between $\overline{Y}$ for a stabilizer state acted upon by $U_{A}\otimes U_{B}$ and that for a Haar random state decays exponentially with the amount of entanglement. In other words, despite the applied unitary being a tensor product of local unitaries on subsystems $A$ and $B$, from the perspective of magic, its effect on magic injection closely resembles that of a genuinely global Haar-random unitary, provided that the initial state is sufficiently entangled. Interestingly, the amount of entanglement necessary for such indistinguishability from Haar random magic is independent of system size. In Fig. 2(b), we numerically simulate this set up using the initial state (2) and averaging over Haar random unitaries $U_{A}$ and $U_{B}$. Again, we find excellent agreement with the prediction of Eq. (LABEL:eq:UaUb), with the subleading corrections negligible for moderate subsystem sizes and entanglement. Moreover, the average $\overline{Y}$ saturates to that of a Haar random state already at $E\approx 3$, which is independent of system size.
Tripartite entanglement.- The above results for bipartite systems can be naturally extended to tripartite systems. Consider a stabilizer state shared among three parties $A$, $B$ and $C$. Since any tripartite stabilizer state can be transformed via local unitaries acting on $A$, $B$, and $C$ into a collection of GHZ state, Bell pair states, and single-qubit states [48], we can, without loss of generality, consider initial states as depicted in Fig. 1(e). Specifically, the initial state consists of $g$ GHZ states shared among $A$, $B$, and $C$; $b_{AB}$, $b_{AC}$, and $b_{BC}$ Bell pairs shared between subsystems $AB$, $AC$, and $BC$, respectively, and $f_{A}$, $f_{B}$, $f_{C}$ single-qubit states in each individual subsystem. Since magic injection via unitary operations on any single subsystem in this setting mirrors the bipartite case discussed earlier, we first consider unitaries of the form $U=U_{A}\otimes U_{B}$ [Fig. 1(c)]. In the limit where $|A|\gg 1$ and $|B|\gg 1$, we find that the average $\overline{Y}$ after applying $U$ is given by [45]
$\overline{Y}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}Y\left(U_{A}U_{B}|\psi\rangle\langle\psi|U_{A}^{\dagger}U_{B}^{\dagger}\right)$ (5)
$=$ $4\cdot{2}^{-|A|-|B|-g-{{b}_{AC}}-{{b}_{BC}}}\left[1+3\cdot{2}^{-2{{b}_{AB}}-g}\right],$
where subleading corrections are of order $O(2^{-|A|},2^{-|B|})$. Eq. (5) can be understood as follows. The average value $\overline{Y}$ is once again exponentially suppressed in both the total size of the subregion where the unitary is applied: $|A|+|B|$, and the amount of entanglement between subregion $AB$ and its complement: $g+b_{AC}+b_{BC}$. Hence the physical interpretation of the prefactor in Eq. (5) is identical to that of Eq. (1). The discrepancy between these two cases lies in the additional factor $3\cdot{2}^{-2{{b}_{AB}}-g}$, which becomes negligible once $A$ and $B$ are sufficiently entangled. This is precisely what we have seen in Eq. (LABEL:eq:UaUb): the effect of a factorized unitary $U_{A}\otimes U_{B}$ becomes essentially indistinguishable from that of a global unitary from the standpoint of magic, once $A$ and $B$ are sufficiently entangled. Thus, Eq. (5) encapsulates the combined effects of Eqs. (1) and (LABEL:eq:UaUb) in the tripartite setting. In the SM [45], we further consider a factorized unitary acting on all three subregions: $U=U_{A}\otimes U_{B}\otimes U_{C}$, where we confirm that in this case the effect of a factorized Haar random unitary on magic injection again becomes indistinguishable from that of a global Haar random unitary, provided that the subregions are sufficiently entangled with one another.
@@
Non-stabilizer entanglement.- So far, we have focused on the propagation of magic on stabilizer states. However, realistic dynamics almost always contains magic in the initial state. Does non-stabilizerness of this kind impose any road blocks on the magic highway? Our findings strongly suggest that, to leading order, the answer is no. Again consider an entangled initial state on $AB$ where $U_{A}$ is applied to inject magic. Since any initial magic local to $A$ can be absorbed via the averaging of $U_{A}$, let us start with a state where non-stabilizerness only lives in the quantum correlation. Such a state is known to have non-local magic, which is non-trivial as long as the entanglement spectrum is non-flat [31]. The simplest example of such a state consists of $k$ imperfect Bell pairs shared between $A$ and $B$: $|\psi\rangle=|0\rangle^{\otimes f_{A}}|\phi_{\theta}\rangle_{AB}^{\otimes k}$, where $|\phi_{\theta}\rangle=\cos\theta|00\rangle+\sin\theta|11\rangle$. Since $k$ copies of $|\phi_{\theta}\rangle$ each with entanglement $S$ can be unitarily distilled into $E=Sk-O(\sqrt{k})$ number of perfect Bell states tensoring a magic state on the $O(\sqrt{k})$ remaining qubits to good approximation [49], our previous analysis with stabilizer initial state again applies to leading order. That is, magic injection by $U_{A}$ will increase in the same way as one increases the total entanglement $E$. Interestingly, this simple dependence also appears to extend to other states with non-local magic. For example, initial states with non-local magic can be parametrized in the form $|\psi\rangle=|0\rangle^{\otimes f_{A}}\sum_{i=1}^{2^{k}}\lambda_{i}|i\rangle_{A}\otimes|i\rangle_{B}$, where $|i\rangle$ denotes a computational basis state of $k$ qubits. We choose four arbitrary entanglement spectra: (i) $\lambda_{i}\propto e^{-i/2^{k}}$; (ii) $\lambda_{i}\propto i$; (iii) $\lambda_{i}\propto i^{2}$; (iv) $\lambda_{i}\propto i^{3}$ and examine the magic spread numerically. Magic propagation is again displaying an $E$ dependence similar to having a stabilizer state or imperfect Bell states as the initial state [Fig. 2.(f)].
@@
Nonetheless, more care is needed going beyond leading order. It is clear that there is non-stabilizerness dependence even in the simplest example with imperfect Bell states at the order of $O(\sqrt{k})$. Furthermore, because initial states can contain magic, the remaining magic capacity for further injection is reduced, i.e., the change of total magic per magic state injection is smaller compared to stabilizer entanglement. We numerically calculate both the magic already existing in the initial states $\overline{Y}_{\text{ini}}$ and the change after the action of $U_{A}$: $\overline{Y}_{\text{final}}/\overline{Y}_{\text{ini}}$. As shown in Fig. 2(d,e), $\overline{Y}_{\rm ini}$ itself in general decays with the increase of entanglement. Nonetheless, we find that the ratio $\overline{Y}_{\text{final}}/\overline{Y}_{\text{ini}}$ also decays exponentially with the amount of entanglement. For the same $E$, the change in global magic becomes smaller as the initial state becomes more magical.
Discussion.- In this work, we demonstrate that stabilizer entanglement functions as a highway that facilitates the spread of locally injected magic throughout the entire system. In particular, we consider an analytically tractable setup where a local Haar random unitary acts on a subregion of a non-magical but potentially highly entangled state, thereby separating the process of magic injection from entanglement generation. We find that the average magic, as quantified by the linear stabilizer entropy, is enhanced by stabilizer entanglement. We further generalize our results to tripartite entanglement and non-stabilizer entanglement. A few interesting open questions remain. Our results for imperfect Bell-pair states suggest that the average amount of injected magic decreases as the amount of magic already existing in the target state increases. This implies that there may be a notion of intrinsic magic capacity of a given state that depends on both its entanglement and magic. Addressing this question would require a more systematic extension of our results to non-stabilizer states, which we leave for future work. Secondly, once we replace the Haar random $U_{A}$ with a more structured brickwork circuit, the question of what is the minimal circuit depth required for the effect of magic injection to coincide with that of a Haar random $U_{A}$ is worth exploring. The question is perhaps intimately related to the complexity growth of local random unitary circuits with circuit depth.
@@
Acknowledgement.- We thank Cheng Wang, Yingfei Gu, and Yuzhen Zhang for helpful discussions, and Xhek Turkeshi for useful comment on the manuscript. We especially thank Daniele Iannotti, Gianluca Esposito, Lorenzo Campos Venuti, and Alioscia Hamma for sharing their note on obtaining the exact expression of average stabilizer purity at a fixed amount of entanglement. This work is supported by Grant No. 12375027 from the National Natural Science Foundation of China (Z.-C. Y.). Numerical simulations were performed on the High-performance Computing Platform of Peking University.
Supplemental Material for “Stabilizer entanglement as a magic highway”
Appendix A Analytical method
@@
In this section, we present the analytical method used to derive the main results quoted in the main text. In Section A.1, we provide a brief introduction to a diagrammatic approach for computing polynomials of matrix elements of unitary matrices averaged over the Haar ensemble. Section A.2 introduces a key quantity used throughout our analysis and outlines approximations made to simplify both the calculation and the final results. Finally, in Section A.3, we detail the step-by-step derivation of the results presented in the main text. Exact results of all setups considered in the main text (i.e. without any approximation such as $|A|\gg 1$) can also be obtained using an alternative analytical method, as provided in Appendix E. This alternative approach, while more compatible with symbolic computations, offers less physical intuition.
A.1 A diagrammatic approach
@@
A.2 A useful quantity and its approximations
In this subsection, we define a quantity that will be used intensively in evaluating $\overline{Y}$. We will calculate the value of this quantity in its most general form, to leading order in $2^{|A|+E}$, where $E$ is the entanglement entropy. Let $\mathbb{S}_{A}$ denote an $|A|$-qubit stabilizer group whose size is $2^{f_{A}}$ with $f_{A}<|A|$; then let $u_{i},i=1,2,3,4$ denote four Pauli strings supported on $A$ (where some of them can be identical) that commute with every element in $\mathbb{S}_{A}$ (i.e. $u_{i}$ is in the normalizer group of $\mathbb{S}_{A}$). Notice that we allow $u_{i}=I_{A}$, but otherwise $u_{i}\notin\mathbb{S}_{A}$. Since we assume $f_{A}<|A|$, one can always find a nontrivial $u_{i}$. And because $u_{i}\mathbb{S}_{A}=u_{i}g\mathbb{S}_{A},\forall g\in\mathbb{S}_{A}$, we can choose one specific $u_{i}$. Then we define the following function:

### v3

@@ Abstract
where $N=|A|+|B|$ is the total number of qubits. First of all, Eq. (LABEL:eq:UaUb) is consistent with Eq. (2) and our teleportation picture upon taking $|A|=|B|=E$. Furthermore, notice that the average value of $\overline{Y^{\rm lin}}$ for Haar random states is given by $1-4\cdot 2^{-N}$. Eq. (LABEL:eq:UaUb) shows that the gap between $\overline{Y^{\rm lin}}$ for a stabilizer state acted upon by $U_{A}\otimes U_{B}$ and that for a Haar random state decays exponentially with the amount of entanglement. Therefore, despite the applied unitary being a tensor product of local unitaries on subsystems $A$ and $B$, from the perspective of magic, its effect on magic injection closely resembles that of a genuinely global Haar-random unitary on $AB$, provided that the initial state is sufficiently entangled. Interestingly, the amount of entanglement necessary for such indistinguishability from Haar random magic is independent of system size. Fig. 2(b) shows numerical simulations of Haar-averaged $U_{A}$ and $U_{B}$, which again closely match Eq. (LABEL:eq:UaUb) with subleading corrections already negligible at moderate $|A|$ and $E$. Remarkably, $\overline{Y^{\rm lin}}$ saturates to the Haar-random value once $E\approx 3$, independent of system size.
In SM, we generalize Eqs. (2) and (LABEL:eq:UaUb) to tripartite states, where similar results are obtained [49].
T count and unitary SRE.- The preceding results establish that entanglement enhances magic injection, but they also raise a natural question: what sets the ultimate limit of this enhancement? From a resource-theoretic perspective, all non-Clifford resources originate from the applied unitary $U_{A}$, so the total injected magic must be upper bounded by the non-Clifford content of $U_{A}$ itself. We begin by introducing the unitary stabilizer Rényi entropy of a given unitary, which provides a direct bridge between our previous results and the non-Clifford resources contained in $U$.
Definition.
Let $U$ be an $N$-qubit unitary, then the $\alpha$-th unitary stabilizer Rényi entropy of $U$ is defined as
[…]
where $P_{i}\in P_{N}$.
To gain some intuition on the above quantity, notice that $UP_{i}U^{\dagger}=\sum c_{ij}P_{j}$ for a single Pauli string $P_{i}$, where $c_{ij}=2^{-N}{\rm tr}(P_{j}UP_{i}U^{\dagger})$, and $\sum_{j}|c_{ij}|^{2}=1$. Thus Eq. (6) is essentially the logarithm of $2^{-2N}\sum_{ij}|c_{ij}|^{2\alpha}$, where the additional factor of $2^{-2N}$ is due to the summation over all $P_{i}$. The definition clearly parallels the state SRE, but now for the unitary directly, which basically characterizes the similarity between $U$ and a Clifford unitary for which $c_{ij}=\pm 1$ for one specific $j$.
We prove that this measure establishes a lower bound on the $T$-count, i.e. the minimal number of $T$ gates needed to synthesize a unitary $U$. More precisely,
Theorem 1.
For any $\alpha\geq 0$,
$t(U)\geq v(U)\geq H_{\alpha}(U),$ (7)
where the unitary stabilizer nullity is defined as $v(U):=2N-{\rm log}|s(U)|$, where $s(U):=\{P_{i}\in P_{N}|UP_{i}U^{\dagger}=\pm P_{j}\}$ [51].
If the total amount of magic of the unitary is given by the unitary SRE above, what initial stabilizer state $|\psi\rangle$, if it exists, can be used to extract all the magic hidden in $U_{A}$? The answer is simply a state that is maximally entangled between $A$ and an ancillary system $B$ of the same size.
Theorem 2.
The unitary SRE of $U$ is equal to the state SRE of $U$’s Choi state $|U\rangle:=U_{A}\otimes I_{B}|\rm Bell\rangle_{AB}^{\otimes N}$:
$H_{\alpha}(U)=M_{\alpha}(|U\rangle).$ (8)
Notice that the right hand side of Eq. (8) is precisely the kind of setup we consider in this work, namely, a unitary acts on a stabilizer initial state consisting of $N$ Bell pairs shared between subsystem $A$ and its complement. Explicitly, the linear stabilizer entropy of acting $U$ on subsystem $A$ when the initial stabilizer state has $|A|=E\leq|B|$ is $Y^{\rm lin}=1-\exp(-H_{2}(U))$.
The above theorem also provides an intuitive interpretation of the unitary SRE we introduce — if we unfold a unitary operator into its dual state using the channel-state duality, then the magic of a unitary operator is precisely equal to that of its dual (Choi) state. Detailed proofs of Theorem 1 and 2 are given in the SM [49].
Figure 3: Numerical simulations of magic injection to initial states with non-stabilizer entanglement. (a)&(b): imperfect Bell states; (c)&(d): states with engineered entanglement spectra (see text for a precise description). In (a)&(c), we plot the amount of preexisting magic in the initial states, while in (b)&(d) we plot the amount of injected magic: $\Delta\overline{M_{2}}=\overline{M_{2}^{\rm final}}-\overline{M_{2}^{\rm init}}$.
@@
Non-stabilizer entanglement.- So far, we have focused on initial states with only stabilizer entanglement. However, realistic dynamics almost always contains magic in the initial state. Does entanglement still enhance magic injection when the initial state is non-stabilizer? Our results indicate that, at leading order, the answer is yes. Consider an entangled state on $AB$ with magic injected by $U_{A}$. Because any local non-stabilizerness in $A$ is washed out by averaging over $U_{A}$, it suffices to focus on states where the non-stabilizer content resides entirely in the correlations between $A$ and $B$. Such a state is known to have non-local magic, which is non-trivial as long as the entanglement spectrum is non-flat [37]. The simplest example of such a state consists of $k$ imperfect Bell pairs shared between $A$ and $B$: $|\psi\rangle=|0\rangle^{\otimes f_{A}}|\phi_{\theta}\rangle_{AB}^{\otimes k}$, where $|\phi_{\theta}\rangle=\cos\theta|00\rangle+\sin\theta|11\rangle$. Since $k$ copies of $|\phi_{\theta}\rangle$ each with entanglement $S$ can be unitarily distilled into $E=Sk-O(\sqrt{k})$ number of perfect Bell states tensoring a magic state on the $O(\sqrt{k})$ remaining qubits to good approximation [52], our previous analysis with stabilizer initial state again applies to leading order. That is, magic injection by $U_{A}$ will increase in the same way as one increases the total entanglement $E$. Interestingly, this simple dependence also appears to extend to other states with non-local magic. For example, initial states with non-local magic can be parametrized in the form $|\psi\rangle=|0\rangle^{\otimes f_{A}}\sum_{i=1}^{2^{k}}\lambda_{i}|i\rangle_{A}\otimes|i\rangle_{B}$, where $|i\rangle$ denotes a computational basis state of $k$ qubits. We choose four arbitrary entanglement spectra: (i) $\lambda_{i}\propto e^{-i/2^{k}}$; (ii) $\lambda_{i}\propto i$; (iii) $\lambda_{i}\propto i^{2}$; (iv) $\lambda_{i}\propto i^{3}$ and examine the magic growth numerically.
@@
However, more care is needed going beyond leading order. It is clear that there is non-stabilizerness dependence even in the simplest example with imperfect Bell states at the order of $O(\sqrt{k})$. Furthermore, because initial states can contain magic, the remaining magic capacity for further injection is reduced, i.e., the change of total magic per magic state injection is smaller compared to stabilizer entanglement. We numerically calculate both the preexisting magic in the initial states and the injected magic by $U_{A}$, as quantified by the increase in the second SRE: $\Delta\overline{M_{2}}=\overline{M_{2}^{\rm final}}-\overline{M_{2}^{\rm init}}$. As shown in Fig. 3(a,c), $\overline{Y^{\rm lin}}$ of both classes of initial states grow with entanglement, indicating that on average the initial state contains more magic when more entangled. The injected magic $\Delta{\overline{M_{2}}}$ also increases with the amount of entanglement [Fig. 3(b,d)]. This suggests that even non-stabilizer entanglement enhances the effect of magic injection, albeit at a much smaller rate. Indeed, as seen from Fig. 3, the amount of injected magic $\Delta\overline{M_{2}}$ becomes smaller as the initial state becomes more magical at the same $E$.
Discussion.- In this work, we demonstrate that entanglement enhances the effect of locally injected magic. We showed this in two ways — by locally injecting magic using a Haar random unitary, and by directly computing the upper limit of maximally extractable magic using unitary SRE. The former shows a precise increase of magic as a function of entanglement while the latter proves that the upper limit is attained when the initial state is maximally entangled. We further generalize our results to tripartite entanglement and non-stabilizer entanglement. As any stabilizer state is equivalent to the tensor product of Bell states, GHZ states and unentangled ancilla through local Clifford rotations, our results apply generally to any bipartite and tripartite stabilizer state. On the resource-theoretic front, we define a new computable measure $H_{\alpha}(U)$ for estimating $t(U)$, which leads to a natural interpretation of entanglement-enhanced magic injection as finding an initial stabilizer state that most efficiently extract the magic contained in a given unitary. Furthermore, it does not rely on optimization and can be computed far more efficiently via tensor network methods, unlike unitary or state nullity. This makes $H_{\alpha}$ a more computable unitary magic measure and a powerful estimate of unitary magic.
Several open questions remain. Our analysis of imperfect Bell-pair states suggests that the average amount of injected magic decreases as the target state itself becomes more magical. This points to a possible notion of an intrinsic magic capacity for a quantum state, determined jointly by its entanglement and existing magic. A systematic extension of our results to generic non-stabilizer states would be needed to make this precise. Another direction concerns replacing the Haar-random $U_{A}$ with a structured brickwork circuit: what minimal circuit depth is required for the injected magic to match that of a Haar-random unitary? This question is likely connected to the growth of complexity in local random unitary circuits as a function of depth.
@@
Acknowledgement.- We thank Zi-Wen Liu, Jong-Yeon Lee, Cheng Wang, Yingfei Gu, and Yuzhen Zhang for helpful discussions, and Xhek Turkeshi for useful comments on the manuscript. We especially thank Daniele Iannotti, Gianluca Esposito, Lorenzo Campos Venuti, and Alioscia Hamma for sharing their note on obtaining the exact expression of average stabilizer purity at a fixed amount of entanglement. Z.-C.Y. is supported by Grant No. 12375027 from the National Natural Science Foundation of China. Numerical simulations were performed on the High-performance Computing Platform of Peking University.
Supplemental Material for “Stabilizer Entanglement Enhances Magic Injection”
Appendix A Additional numerical results on the second stabilizer Rényi entropy $\overline{M_{2}}$
In this section, we provide additional numerical results on the Haar-average value of the second stabilizer Rényi entropy $\overline{M_{2}}$, for the setup depicted in Fig. 1(a) of the main text. We find that the the average $\overline{M_{2}}$ is extremely close to ${\rm log}(1-\overline{Y^{\rm lin}})$, see Fig. 4. This indicates that the average $\overline{Y^{\rm lin}}$ is also typical, as sample-to-sample fluctuations are small.
Figure 4: Numerical simulations of $\overline{Y^{\rm lin}}$ and $\overline{M_{2}}$ for the setup depicted in Fig.1(a) of the main text. $1-\overline{Y^{\rm lin}}$ is shown as upper triangles and $2^{-\overline{M_{2}}}$ is shown as lower triangles.
Appendix B Tripartite entanglement
In this section, we extend the main results presented in the main text to tripartite entanglement. Consider a stabilizer state shared among three parts $A$, $B$ and $C$. Since any tripartite stabilizer state can be transformed via local unitaries acting on $A$, $B$, and $C$ into a collection of GHZ states, Bell states, and single-qubit states [50], we can, without loss of generality, consider initial states as depicted in Fig. 5(a). Specifically, the initial state consists of $g$ GHZ states shared among $A$, $B$, and $C$; $b_{AB}$, $b_{AC}$, and $b_{BC}$ Bell states shared between subsystems $AB$, ACAC, and $BC$, respectively, and $f_{A}$, $f_{B}$, $f_{C}$ single-qubit states in each individual subsystem. Since magic injection via unitary operations on any single subsystem in this setting mirrors the bipartite case discussed earlier, we first consider unitaries of the form $U=U_{A}\otimes U_{B}$ [Fig. 5(b)]. In the limit where $|A|\gg 1$ and $|B|\gg 1$, we find that the average $\overline{Y^{\rm lin}}$ after applying $U$ is given by
$\overline{Y^{\rm lin}}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}Y^{\rm lin}\left(U_{A}U_{B}|\psi\rangle\langle\psi|U_{A}^{\dagger}U_{B}^{\dagger}\right)$ (9)
$=$ $1-4\cdot{2}^{-|A|-|B|-g-{{b}_{AC}}-{{b}_{BC}}}\left[1+3\cdot{2}^{-2{{b}_{AB}}-g}\right],$
where subleading corrections are of order $O(2^{-|A|},2^{-|B|})$. A detailed derivation of Eq. (9) is given in Sec. C.3. Eq. (9) can be understood as follows. The average value $1-\overline{Y^{\rm lin}}$ is once again exponentially suppressed in both the total size of the subregion where the unitary is applied: $|A|+|B|$, and the amount of entanglement between subregion $AB$ and its complement: $g+b_{AC}+b_{BC}$. Hence the physical interpretation of the prefactor in Eq. (9) is identical to that of Eq. (2) in the main text. The discrepancy between these two cases lies in the additional factor $3\cdot{2}^{-2{{b}_{AB}}-g}$, which becomes negligible once $A$ and $B$ are sufficiently entangled. This is precisely what we have seen in Eq. (LABEL:eq:UaUb) of the main text: the effect of a factorized unitary $U_{A}\otimes U_{B}$ becomes essentially indistinguishable from that of a global unitary from the standpoint of magic, once $A$ and $B$ are sufficiently entangled. Thus, Eq. (9) encapsulates the combined effects of Eqs. (2) and  (LABEL:eq:UaUb) of the main text in the tripartite setting. In Sec. E, we further consider a factorized unitary acting on all three subregions: $U=U_{A}\otimes U_{B}\otimes U_{C}$, where we confirm that in this case the effect of a factorized Haar random unitary on magic injection again becomes indistinguishable from that of a global Haar random unitary, provided that the subregions are sufficiently entangled with one another.
Figure 5: (a) Schematic of the initial stabilizer state used for tripartite systems. Notice that any tripartite stabilizer state can be brought into this form via local Clifford unitaries acting on subsystems $A$, $B$ and $C$. (b) Schematic of the setup for $U_{A}\otimes U_{B}$ acting on tripartite systems. (c) Numerical results for $U_{A}\otimes U_{B}$ acting on tripartite systems. The dashed lines represent the quantity shown in Eq. (9) and the squares represent the exact result Eq. (114).
Appendix C Analytical method
@@
In this section, we present the analytical method used to derive the main results quoted in the main text. In Section C.1, we provide a brief introduction to a diagrammatic approach for computing polynomials of matrix elements of unitary matrices averaged over the Haar ensemble. Section C.2 introduces a key quantity used throughout our analysis and outlines approximations made to simplify both the calculation and the final results. Finally, in Section C.3, we detail the step-by-step derivation of the results presented in the main text. Exact results of all setups considered in the main text (i.e. without any approximation such as $|A|\gg 1$) can also be obtained using an alternative analytical method, as provided in Appendix H. This alternative approach, while more compatible with symbolic computations, offers less physical intuition.
C.1 A diagrammatic approach
@@
C.2 A useful quantity and its approximations
In this subsection, we define a quantity that will be used intensively in evaluating $\overline{Y^{\rm lin}}$. We will calculate the value of this quantity in its most general form, to leading order in $2^{|A|+E}$, where $E$ is the entanglement entropy. Let $\mathbb{S}_{A}$ denote an $|A|$-qubit stabilizer group whose size is $2^{f_{A}}$ with $f_{A}<|A|$; then let $u_{i},i=1,2,3,4$ denote four Pauli strings supported on $A$ (where some of them can be identical) that commute with every element in $\mathbb{S}_{A}$ (i.e. $u_{i}$ is in the normalizer group of $\mathbb{S}_{A}$). Notice that we allow $u_{i}=I_{A}$, but otherwise $u_{i}\notin\mathbb{S}_{A}$. Since we assume $f_{A}<|A|$, one can always find a nontrivial $u_{i}$. And because $u_{i}\mathbb{S}_{A}=u_{i}g\mathbb{S}_{A},\forall g\in\mathbb{S}_{A}$, we can choose one specific $u_{i}$. Then we define the following function: